OpenStudy (anonymous):

Weights of male mountain lions follow the normal distribution with a median of 150 pounds and an interquartile range of 8.2 pounds. Find the 75th percentile of the weights.

OpenStudy (kirbykirby):

In a normal distribution, the median = mean. Also, The mean = 50th percentile ("Q2") The interquartile range = Q3 - Q1 = 75th percentile - 25th percentile. Now, since the normal distribution is symmetric, half your data is below the mean, and half is above the mean. So, you proportionately find 8.2/2 = 4.1 pounds being Q3 - Q2 , and Q2 - Q1 Schematically: